OpenStudy (lilsis76):

A 100.0 mL sample of water is heatedto its Boiling Point. How much heat (in kj) is required to vaporize it? (Assume Density of 1.00 g/mL). Please Help. Word problems freak me out, and I really dont understand the concept of conversion if someone can help me understand and answer this problem please. THank you.

OpenStudy (aaronq):

if the water is already at 100 celsius, then you only need to multiply the mass or moles (depending on what the \(\Delta H_{vap}\) given is in units of. q=\(\Delta H_{vap}\) *m orq=\(\Delta H_{vap}\) *n

OpenStudy (lilsis76):

@aaronq but I dont get it, so OH! wait boiling point of water is at 100 degree Celcius. so I multiply by H=1.008 and O=16 ?

OpenStudy (aaronq):

hm no you're confusing things. Boiling point and molar masses aren't related like that. DId they give you a value for \(\Delta H_{vap}\)?

OpenStudy (lilsis76):

all i have from the quesiton is:.... 100.0mL sample of water heated to its boiling point. ?___in kilojoules. Density of 1.00 g/mL. And in the book it has...... liquid water, T=20 deg. Celcious, Density is .998 g/cm cubed at 1 atm on page 456

OpenStudy (lilsis76):

oh, that was from the 3 states of water in chap. 11

OpenStudy (lilsis76):

FOUND IT!

OpenStudy (lilsis76):

okay, page 472. Heats of vaporization at their boiling points and at 25 deg. C. Waters normal boiling point is 100 deg cel. heat of vap(kJ/mol) at 25 deg C is 44.0

OpenStudy (lilsis76):

is that what I need?

OpenStudy (aaronq):

You need it to be at 100 celsius, it's \(\Delta H_{vap}\)= 40.7 kJ/mol (usually in a test they give you this, if not memorize it)

OpenStudy (aaronq):

so all you need to do is to convert the water given to grams, then to moles. \(\rho=\dfrac{m}{V}\); where, \(\rho\)=density then, \(n=\dfrac{m}{M}\) then, q=\(\Delta H_{vap}*n\) make sure they units are the same.

OpenStudy (lilsis76):

okay, sorry could you watch me on here to see if i do it right. okay so D=m/vol. Density=\[\frac{ 18g }{ 100mL }\]

OpenStudy (lilsis76):

is that right? @aaronq

OpenStudy (aaronq):

where'd you get 18 g from? what you're solving is for mass

OpenStudy (aaronq):

ohh, m is not M, they're not asking for molar mass 1.0 g/mL =\(\dfrac{m}{100 mL}\)

OpenStudy (lilsis76):

ya, I added he H2O together. = 18

OpenStudy (aaronq):

that's molar mass (M) not mass (m)

OpenStudy (aaronq):

yeah, little details, like capitals, make a big difference in chemistry.

OpenStudy (lilsis76):

oh... :? then im completely lost

OpenStudy (aaronq):

rearrange: \(m=\rho*V\) m=1 g/mL * 100 mL = 100 g now, convert to moles (use the molar mass this time)

OpenStudy (lilsis76):

okay so....here hold on

OpenStudy (lilsis76):

okay so....here hold on \[d=\frac{ m }{ v } \]

OpenStudy (aaronq):

volume should be a capital "V", lower case v means velocity

OpenStudy (lilsis76):

okay D=\[\frac{ m }{ V }\]

OpenStudy (aaronq):

haha density should be either "rho" (\(\rho\)) or lower case d :P, but don't worry about it THIS time.

OpenStudy (lilsis76):

now i .....well okay, we have the mass that is the 100grams, the volume is going to be the... 100 mL-->Liters right? so we divide by 1000 giving me 0.1 L

OpenStudy (aaronq):

hmm we already used that equation, remember? we found mass using m=\(\rho*V\)= 100g now we need to find moles: \(n=\dfrac{m}{M}\)

OpenStudy (lilsis76):

okay so n= mass being the 100g / M-Molar mass? being 18 grams 100g/18g =5.56?

OpenStudy (lilsis76):

@aaronq did i do it right this time?

OpenStudy (aaronq):

yep thats right. molar mass is in units of "g/mol" though.

OpenStudy (aaronq):

as in "18 g is equal to 1 mole of water"

OpenStudy (lilsis76):

oh okay hold on let me write that down

OpenStudy (lilsis76):

so then the answer @aaronq is 5.56 moles of kJ is needed to heat the sample?

OpenStudy (aaronq):

nope, you can't express energy in terms of moles :P q=n*\(\Delta H_{vap}\) and you're done

OpenStudy (aaronq):

q=heat

OpenStudy (lilsis76):

lol let me try

OpenStudy (lilsis76):

i cant do it..............:(

OpenStudy (aaronq):

lol why not? q=n*ΔHvap =5.56 moles * 40.7 kJ/mol = 226.292 kJ

OpenStudy (lilsis76):

because im crying thats why. haha its sad

OpenStudy (aaronq):

lol this shouldn't be that bad.. :P are you planning to study chemistry further?

OpenStudy (lilsis76):

yes. I took chemistry 101 and passed with a C, took chem 102 but was in a car accident. returned to school and had to take a failing grade in order to retake it, I took it again and I failed. I took it again and I withdrew cuz I realized I was rushing myself to get back into school of 2 years not taking chemistry. and Now I kind of know it but I dont get conversions and I have no idea what all the equaitons and conversion factors are. do you know of a website that has them?? I really wanna be a biology major to be a doctor

OpenStudy (lilsis76):

sorry about my sad life story but Im so stressed

OpenStudy (aaronq):

sorry to hear that :S um a website with conversion factors? no particularly, but if you google whatever specific conversion you want, it'll usually pop up in yahoo questions or something.

OpenStudy (lilsis76):

its okay, thank you. And yes ive done that. but I think I might need it for test taking :/ I dunno. ugh...thank you for helping me out.

OpenStudy (lilsis76):

Im going to close this problem. would you mind helping me with one more problem, of many homework quesitons that I dont get? @aaronq

OpenStudy (aaronq):

You should make a list of all the equation as you go along. and yeah, no problem. I have to go in a few minutes so please post it fast :P

OpenStudy (lilsis76):

okay. and I have so many questions but right now its why cant i stop sniffling. haha :')

OpenStudy (aaronq):

haha dont stress yourself out, it doesn't help! :P calm and collected is where it's at.

OpenStudy (lilsis76):

i know, i know. Im trying so hard. BLAH! haha